# Propensity score weighting across counterfactual worlds: longitudinal effects under positivity violations

### v1

@@ 1 Introduction
More successful approaches for addressing positivity violations in longitudinal data have stemmed from “stochastic longitudinal modified treatment policies” (S-LMTPs). First introduced in Robins et al. (2004) and Stock (1989), S-LMTPs have subsequently evolved through various contributions (van der Laan and Petersen, 2007; Haneuse and Rotnitzky, 2013; Díaz and van der Laan, 2012; Kennedy, 2019; Díaz et al., 2023; Richardson and Robins, 2013; Young et al., 2014). Among these, a recent development for weighting is the “flip” intervention, which flips subjects toward a target treatment regime using weights constructed from non-baseline covariates (McClean et al., 2025). This construction allows for dynamic weighting and trimming on non-baseline covariates, making it possible to avoid positivity violations at each timepoint. However, flip interventions face an inherent limitation: because they are “single-world” interventions (in the sense of Richardson and Robins (2013)) that sequentially adapt to prior counterfactual interventions, contrasts between potential outcomes under two sequences of flip interventions do not isolate the difference in mean potential outcomes under the two static regimes of interest. The dynamic nature of flip interventions means that the intervention at timepoint $t$ affects not only the final outcome but also subsequent covariates and treatments, which in turn affects future flip interventions. Consequently, differences in flip effects can reflect the intervention’s impact on the entire treatment process rather than just the ultimate outcome, with the corresponding implications for interpretation.
@@ 1.1 Our contributions
$\psi(\overline{a}_{T},\overline{a}_{T}^{\prime}):=\mathbb{E}\left(\left\{Y(\overline{a}_{T})-Y(\overline{a}_{T}^{\prime})\right\}\prod_{t=1}^{T}w_{t}\left[p_{t}\{\overline{X}_{t}(\overline{a}_{t-1})\}\right]w_{t}^{\prime}\left[p_{t}^{\prime}\{\overline{X}_{t}(\overline{a}_{t-1}^{\prime}\}\right]\right),$ (1)
where $p_{t}\{\overline{X}_{t}(\overline{a}_{t-1})\}=\mathbb{P}\{A_{t}(\overline{a}_{t-1})=a_{t}\mid\overline{X}_{t-1}(\overline{a}_{t-1})\}$ and $p_{t}^{\prime}\{\overline{X}_{t}(\overline{a}_{t-1}^{\prime}\}=\mathbb{P}\{A_{t}(\overline{a}_{t-1}^{\prime})\mid X_{t}(\overline{a}_{t-1}^{\prime})\}$ are the natural propensity scores under interventions $\overline{a}_{t-1}$ and $\overline{a}_{t-1}^{\prime}$, respectively, and $w_{t},w_{t}^{\prime}$ are weighting functions.
@@ 1.3 Mathematical notation
[…]
@@ 2 Setup for longitudinal data
Here, $\Big{\{}\big{\{}U_{X,t},U_{A,t}:t\in\{1,\ldots,T\}\big{\}},U_{Y}\Big{\}}$ is a vector of exogenous variables. Subsequently, we’ll define restrictions on their joint distribution that facilitate identification of causal effects. We will define the effects in terms of hypothetical interventions in which equation $A_{t}=f_{A,t}(H_{t},U_{A,t})$ is removed from the structural model and the exposure is assigned as a new treatment $a_{t}$. An intervention that sets exposures up to time $t-1$ to $\overline{a}_{t-1}$ generates counterfactual covariates $X_{t}(\overline{a}_{t-1})=f_{X,t}\big{\{}a_{t-1},H_{t-1}(\overline{a}_{t-2}),U_{X,t}\big{\}}$, where the counterfactual treatment and covariate history is defined recursively as $H_{t}(\overline{a}_{t-1})=\{X_{t}(\overline{a}_{t-1}),a_{t-1},H_{t-1}(\overline{a}_{t-2})\}$. The variable $A_{t}(\overline{a}_{t-1})$ is called the natural treatment value (Richardson and Robins, 2013; Young et al., 2014), and represents the possibly counterfactual value of the treatment that would have been observed at time $t$ under an intervention carried out up to time $t-1$ but discontinued thereafter. Similarly, we refer to $X_{t}(\overline{a}_{t-1})$ as the natural covariate value. Moreover, we define the counterfactual covariate history at time $t$ under intervention $\overline{a}_{t-1}$ as $\overline{X}_{t}(\overline{a}_{t-1})=\big{\{}X_{s}(\overline{a}_{s-1})\big{\}}_{s=1}^{t}$. Finally, an intervention in which all treatment variables up to $t=T$ are intervened on generates a counterfactual outcome $Y\left(\overline{a}_{T}\right)=f_{Y}\big{\{}a_{T},H_{T}(\overline{a}_{T-1}),U_{Y}\big{\}}$.
@@ 2.1 Causal assumptions
Assumption 1 is satisfied if common causes of treatment $A_{t}$ and future covariates and treatments are measured. This assumption is similar to that required by Richardson and Robins (2013) (cf. Theorem 31). Finally, note that we do not require the typical positivity assumption, which says that time-varying propensity scores are bounded away from zero and one ($10<\mathbb{P}(A_{t}=1\mid H_{t})<$ 1), because we will consider longitudinal weights which adapt to positivity violations.
@@ 3 Cumulative cross-world weighted treatment effects
$\mathbb{E}\left(\left\{Y(\overline{a}_{T})-Y(\overline{a}_{T}^{\prime})\right\}\prod_{t=1}^{T}w_{t}\left[p_{t}\{\overline{X}_{t}(\overline{a}_{t-1})\}\right]w_{t}^{\prime}\left[p_{t}^{\prime}\{\overline{X}_{t}(\overline{a}_{t-1}^{\prime}\}\right]\right).$ (1)
@@
  • $p_{t}\{\overline{X}_{t}(\overline{a}_{t-1})\}=\mathbb{P}\left\{A_{t}(\overline{a}_{t-1})=a_{t}\mid\overline{X}_{t}(\overline{a}_{t-1})\right\}$ and $p_{t}^{\prime}\{\overline{X}_{t}(\overline{a}_{t-1})\}=\mathbb{P}\left\{A_{t}(\overline{a}_{t-1}^{\prime})=a_{t}^{\prime}\mid\overline{X}_{t}(\overline{a}_{t-1}^{\prime})\right\}$ are the natural propensity scores under regimes $\overline{a}_{t-1}$ and $\overline{a}_{t-1}^{\prime}$, respectively; and
@@
$w_{t}^{\prime}\left[p_{t}^{\prime}\{\overline{X}_{t}(\overline{a}_{t-1})\}\right]$ $=\mathbbold{1}\big{[}p_{t}^{\prime}\{\overline{X}_{t}(\overline{a}_{t-1})\}>\varepsilon^{\prime}\big{]},$
@@
Weighting towards $\overline{a}_{T}^{\prime}$ only $1$ $p_{t}^{\prime}$
@@ 3.1 Benefits and drawbacks
This result shows that the cumulative cross-world weighted effect is zero whenever the two potential outcomes under consideration are almost surely equal. This is a useful result. In particular, alternative proposals, like flip effects, do not satisfy it (cf, McClean et al. (2025), Section 4.2). As discussed in McClean et al. (2025), because flip interventions affect subsequent treatments and outcomes as well as the ultimate outcome of interest, the difference in mean potential outcomes under two flip interventions will typically fail to isolate the difference in the two potential outcomes of interest.
@@
where $\overline{X}_{t,i}(\overline{a}_{t-1}),\overline{X}_{t,i}(\overline{a}_{t-1}^{\prime})$ are the natural covariate and treatment history for unit $i$. Supposing one can identify these weights from the observed data (see, Section 4), it is possible to assign a weight to each individual in the study and examine summary statistics of the weights.
@@ 5.3 Doubly robust-style estimator and convergence guarantees
$\widehat{\psi}(\overline{a}_{T}):=\mathbb{P}_{n}\{\widehat{\varphi}(Z)\}\text{and }\widehat{\sigma}^{2}:=\mathbb{P}_{n}\left[\{\widehat{\varphi}(Z)-\widehat{\psi}(\overline{a}_{T})\}^{2}\right].$
@@ 6.3 Results
Figure 1 displays the distributions of the propensity scores under each of the targeted regimes. Notably, some probabilities are near zero. In the first row of Figure 1, propensity scores near zero indicate that certain workers would have close to zero probability of being in a union if they were in a union previously. In the second row, this indicates that certain workers would have almost probability one of being in a union if they were not in a union previously. This is a reasonable scenario for considering weighting to adapt to positivity violations. Table 2 presents the distribution of the weights and the effective sample size at each time point. The effective sample size at each time point was calculated as the squared sum of cumulative weights up to that time point divided by the sum of squares of cumulative weights. Together, these results show that the weights were not extreme. Table 2 also provides evidence on the common support assumption, in Assumption 2. It shows that the covariate density ratios for timepoints 2 to 4 are are not that extreme, suggesting that the common support assumption may hold for all subjects.
@@
Figure 1: Cumulative cross-world weights.
@@ A.1 Lemma 1
=$=\frac{\mathbb{P}\{A_{t}(\overline{a}_{t-1})=a_{t},X_{t}(\overline{a}_{t-1}),\dots,X_{2}\mid A_{1}=a_{1}X_{1}\}}{d\mathbb{P}\{X_{t}(\overline{a}_{t-1}),\dots,X_{2}\mid A_{1}=a_{1},X_{1}\}}$
@@
where the first line follows by factoring the conditional probability and the second line by strong sequential randomization, the positivity assumption, that $\pi_{1}(X_{1})>0$, so that the conditioning set is well-defined, and by consistency, so that $X_{2}(a_{1})=X_{2}$, conditional on $A_{1}=a_{1},X_{1}$. We leave the other left-hand side variables as their natural values for ease of notation, although one could write $X_{3}(a_{2},A_{1})$ and so on, instead of what we have written. The third line follows by factoring the conditional probabilities again, and the fourth line follows by strong sequential randomization, positivity, and consistency. This argument can be repeated $t-3$ more times to yield
@@ A.2 Theorem 1
$=\mathbb{E}\left\{\mathbb{E}\left(\mathbb{E}\left[Y(\overline{a}_{T})\prod_{t=3}^{T}w_{t}\big{[}\mathbb{P}\{A_{t}(\overline{a}_{t-1})=a_{t}\mid\overline{X}_{t}(\overline{a}_{t-1})\}\big{]}w_{t}^{\prime}\big{[}\mathbb{P}\{A_{t}(\overline{a}_{t-1}^{\prime})=a_{t}\mid\overline{X}_{t}(\overline{a}_{t-1}^{\prime})\}\big{]}\mid X_{2},A_{1}=a_{1},X_{1}\right]w_{2}(\pi_{2})w_{2}^{\prime}(\pi_{2}^{\prime})\mid X_{1}\right)w_{1}(\pi_{1})w_{1}^{\prime}(\pi_{1}^{\prime})\right\}.$
@@ Appendix B Proofs for Section 5
[…]
@@
where the first line follows by repeating, from $s=t-1$ to $s=1$, iterated expectations on $\overline{X}_{s},\overline{A}_{s-1}=\overline{a}_{s-1}$ and then iterated expectations on $\overline{X}_{s-1},\overline{A}_{s-1}=\overline{a}_{s-1}^{\prime}$ and canceling terms, and the second line follows by repeating, from $s=1$ to $t-1$, iterated expectations on $\overline{X}_{s},\overline{A}_{s}=\overline{a}_{s-1}$.
@@
Second-order Taylor expansions of $w_{t}(\pi)$ and $w_{t}^{\prime}(\pi^{\prime})$ around $w_{t}(\widehat{\pi})$ and $w_{t}^{\prime}(\widehat{\pi}^{\prime})$, respectively, yield
@@
Repeating this argument $t-2$ times, and noting that $\rho_{1}=1$, implies that $\mathbb{V}\{\varphi(Z)\}$ is bounded, which yields the result. ∎

### v2

@@ 1 Introduction
More successful approaches for addressing positivity violations in longitudinal data have stemmed from “stochastic longitudinal modified treatment policies” (S-LMTPs). First introduced in Robins et al. (2004) and Stock (1989), S-LMTPs have subsequently evolved through various contributions (van der Laan and Petersen, 2007; Haneuse and Rotnitzky, 2013; D´ıaz and van der Laan, 2012; Kennedy, 2019; D´ıaz et al., 2023; Richardson and Robins, 2013; Young et al., 2014). Among these, a recent development for weighting is the “flip” intervention, which flips subjects toward a target treatment regime using weights constructed from non-baseline covariates (McClean et al., 2025). This construction allows for dynamic weighting and trimming on non-baseline covariates, making it possible to avoid positivity violations at each timepoint. However, flip interventions face an inherent limitation: because they are “single-world” interventions (in the sense of Richardson and Robins (2013)) that sequentially adapt to prior counterfactual interventions, contrasts between potential outcomes under two sequences of flip interventions do not isolate the difference in mean potential outcomes under the two static regimes of interest. The dynamic nature of flip interventions means that the intervention at timepoint $t$ affects not only the final outcome but also subsequent covariates and treatments, which in turn affect future flip interventions. Consequently, differences in flip effects can reflect the intervention’s impact on the entire treatment process rather than just the ultimate outcome, with the corresponding implications for interpretation.
@@ 1.1 Our contributions
$\psi(\overline{a}_{T},\overline{a}_{T}^{\prime}):=\mathbb{E}\left(\left\{Y(\overline{a}_{T})-Y(\overline{a}_{T}^{\prime})\right\}\prod_{t=1}^{T}w_{t}\left[p_{t}\{\overline{X}_{t}(\overline{a}_{t-1})\}\right]w_{t}^{\prime}\left[p_{t}^{\prime}\{\overline{X}_{t}(\overline{a}_{t-1}^{\prime})\}\right]\right),$ (1)
where $p_{t}\{\overline{X}_{t}(\overline{a}_{t-1})\}=\mathbb{P}\{A_{t}(\overline{a}_{t-1})=a_{t}\mid\overline{X}_{t-1}(\overline{a}_{t-1})\}$ and $p_{t}^{\prime}\{\overline{X}_{t}(\overline{a}_{t-1}^{\prime})\}=\mathbb{P}\{A_{t}(\overline{a}_{t-1}^{\prime})=a_{t}^{\prime}\mid X_{t}(\overline{a}_{t-1}^{\prime})\}$ are the natural propensity scores under interventions $\overline{a}_{t-1}$ and $\overline{a}_{t-1}^{\prime}$, respectively, and $w_{t},w_{t}^{\prime}$ are weighting functions.
@@ 1.3 Mathematical notation
For a function $f(Z)$, we use $\lVert f\rVert=\sqrt{\int f(z)^{2}d\mathbb{P}(z)}$ to denote the $L_{2}(\mathbb{P})$ norm, $\lVert f\rVert_{p}=\left(\int f(z)^{p}d\mathbb{P}(z)\right)^{1/p}$ to denote the generic $L_{p}(\mathbb{P})$ norm for $p\geq 1$, $\mathbb{P}(f)=\int_{\mathcal{Z}}f(z)d\mathbb{P}(z)$ to denote the average with respect to the underlying distribution $\mathbb{P}$, and $\mathbb{P}_{n}(f)=\frac{1}{n}\sum_{i=1}^{n}f(Z_{i})$ to denote the empirical average with respect to $n$ observations. In a standard abuse of notation, when $B$ is an event we let $\mathbb{P}(B)$ denote the probability of $B$. We also denote expectation and variance with respect to the underlying distribution by $\mathbb{E}$ and $\mathbb{V}$, respectively. We use $a\wedge b$ for minimum and $a\vee b$ for maximum, and $a\lesssim b$ to mean $a\leq Cb$ for some constant $C$. We use $\rightsquigarrow$ to denote convergence in distribution, and $\stackrel{{\scriptstyle p}}{{\rightarrow}}$ for convergence in probability. Additionally, we use $o_{\mathbb{P}}(\cdot)$ to denote convergence in probability to zero, i.e., if $X_{n}$ is a sequence of random variables then $X_{n}=o_{\mathbb{P}}(r_{n})$ implies $\left|\frac{X_{n}}{r_{n}}\right|\stackrel{{\scriptstyle p}}{{\rightarrow}}0$.
@@ 2 Setup for longitudinal data
Here, $\Big{\{}\big{\{}U_{X,t},U_{A,t}:t\in\{1,\ldots,T\}\big{\}},U_{Y}\Big{\}}$ is a vector of exogenous variables. Subsequently, we’ll define restrictions on their joint distribution that facilitate identification of causal effects. We will define the effects in terms of hypothetical interventions in which equation $A_{t}=f_{A,t}(H_{t},U_{A,t})$ is removed from the structural model and the exposure is assigned as a new treatment $a_{t}$. An intervention that sets exposures up to time t1t-1 to $\overline{a}_{t-1}$ generates counterfactual covariates $X_{t}(\overline{a}_{t-1})=f_{X,t}\big{\{}a_{t-1},H_{t-1}(\overline{a}_{t-2}),U_{X,t}\big{\}}$, where the counterfactual treatment and covariate history is defined recursively as $H_{t}(\overline{a}_{t-1})=\{X_{t}(\overline{a}_{t-1}),a_{t-1},H_{t-1}(\overline{a}_{t-2})\}$. The variable $A_{t}(\overline{a}_{t-1})$ is called the natural treatment value (Richardson and Robins, 2013; Young et al., 2014), and represents the possibly counterfactual value of the treatment that would have been observed at time $t$ under an intervention carried out up to time t1t-1 but discontinued thereafter. Similarly, we refer to $X_{t}(\overline{a}_{t-1})$ as the natural covariate value. Moreover, we define the counterfactual covariate history at time $t$ under intervention $\overline{a}_{t-1}$ as $\overline{X}_{t}(\overline{a}_{t-1})=\big{\{}X_{s}(\overline{a}_{s-1})\big{\}}_{s=1}^{t}$. Finally, an intervention in which all treatment variables up to $t=T$ are intervened on generates a counterfactual outcome $Y\left(\overline{a}_{T}\right)=f_{Y}\big{\{}a_{T},H_{T}(\overline{a}_{T-1}),U_{Y}\big{\}}$.
@@ 2.1 Causal assumptions
Assumption 1 is satisfied if common causes of treatment $A_{t}$ and future covariates and treatments are measured. This assumption is similar to that required by Richardson and Robins (2013) (cf. Theorem 31). Finally, note that we do not require the typical positivity assumption, which says that time-varying propensity scores are bounded away from zero and one ($0<\mathbb{P}(A_{t}=1\mid H_{t})<1$), because we will consider longitudinal weights which adapt to positivity violations.
@@ 3 Cumulative cross-world weighted treatment effects
$\mathbb{E}\left(\left\{Y(\overline{a}_{T})-Y(\overline{a}_{T}^{\prime})\right\}\prod_{t=1}^{T}w_{t}\left[p_{t}\{\overline{X}_{t}(\overline{a}_{t-1})\}\right]w_{t}^{\prime}\left[p_{t}^{\prime}\{\overline{X}_{t}(\overline{a}_{t-1}^{\prime})\}\right]\right).$ (1)
@@
  • $p_{t}\{\overline{X}_{t}(\overline{a}_{t-1})\}=\mathbb{P}\left\{A_{t}(\overline{a}_{t-1})=a_{t}\mid\overline{X}_{t}(\overline{a}_{t-1})\right\}$ and $p_{t}^{\prime}\{\overline{X}_{t}(\overline{a}_{t-1}^{\prime})\}=\mathbb{P}\left\{A_{t}(\overline{a}_{t-1}^{\prime})=a_{t}^{\prime}\mid\overline{X}_{t}(\overline{a}_{t-1}^{\prime})\right\}$ are the natural propensity scores under regimes $\overline{a}_{t-1}$ and $\overline{a}_{t-1}^{\prime}$, respectively; and
@@
$w_{t}^{\prime}\left[p_{t}^{\prime}\{\overline{X}_{t}(\overline{a}_{t-1}^{\prime})\}\right]$ $=\mathbbold{1}\big{[}p_{t}^{\prime}\{\overline{X}_{t}(\overline{a}_{t-1}^{\prime})\}>\varepsilon^{\prime}\big{]},$
@@
Weighting towards $\overline{a}_{T}^{\prime}$ only 11 $p_{t}^{\prime}$
@@ 3.1 Benefits and drawbacks
This result shows that the cumulative cross-world weighted effect is zero whenever the two potential outcomes under consideration are almost surely equal. This is a useful result. In particular, alternative proposals, like flip effects, do not satisfy it (cf. , McClean et al. (2025), Section 4.2). As discussed in McClean et al. (2025), because flip interventions affect subsequent treatments and outcomes as well as the ultimate outcome of interest, the difference in mean potential outcomes under two flip interventions will typically fail to isolate the difference in the two potential outcomes of interest.
@@
where $\overline{X}_{t,i}(\overline{a}_{t-1}),\overline{X}_{t,i}(\overline{a}_{t-1}^{\prime})$ are the natural covariate and treatment history for unit $i$. Supposing one can identify these weights from the observed data (see Section 4), it is possible to assign a weight to each individual in the study and examine summary statistics of the weights.
@@ 5.3 Doubly robust-style estimator and convergence guarantees
[…]
@@ 6.3 Results
Figure 1 displays the distributions of the propensity scores under each of the targeted regimes. Notably, some probabilities are near zero. In the first row of Figure 1, propensity scores near zero indicate that certain workers would have close to zero probability of being in a union if they were in a union previously. In the second row, this indicates that certain workers would have almost probability one of being in a union if they were not in a union previously. This is a reasonable scenario for considering weighting to adapt to positivity violations. Table 2 presents the distribution of the weights and the effective sample size at each timepoint. The effective sample size at each timepoint was calculated as the squared sum of cumulative weights up to that timepoint divided by the sum of squares of cumulative weights. Together, these results show that the weights were not extreme. Table 2 also provides evidence on the common support assumption, in Assumption 2. It shows that the covariate density ratios for timepoints 2 to 4 are not that extreme, suggesting that the common support assumption may hold for all subjects.
@@
Figure 1: Propensity score distribution.
@@ A.1 Lemma 1
$=\frac{\mathbb{P}\{A_{t}(\overline{a}_{t-1})=a_{t},X_{t}(\overline{a}_{t-1}),\dots,X_{2}\mid A_{1}=a_{1}X_{1}\}}{d\mathbb{P}\{X_{t}(\overline{a}_{t-1}),\dots,X_{2}\mid A_{1}=a_{1},X_{1}\}}$
@@
where the first line follows by factoring the conditional probability and the second line by strong sequential randomization, the positivity assumption, that $\pi_{1}(X_{1})>0$, so that the conditioning set is well-defined, and by consistency, so that $X_{2}(a_{1})=X_{2}$, conditional on $A_{1}=a_{1},X_{1}$. We leave the other left-hand side variables as their natural values for ease of notation, although one could write $X_{3}(a_{2},A_{1})$ and so on, instead of what we have written. The third line follows by factoring the conditional probabilities again, and the fourth line follows by strong sequential randomization, positivity, and consistency. This argument can be repeated t3t-3 more times to yield
@@ A.2 Theorem 1
$\hskip-54.2025pt=\mathbb{E}\left\{\mathbb{E}\left(\mathbb{E}\left[Y(\overline{a}_{T})\prod_{t=3}^{T}w_{t}\big{[}\mathbb{P}\{A_{t}(\overline{a}_{t-1})=a_{t}\mid\overline{X}_{t}(\overline{a}_{t-1})\}\big{]}w_{t}^{\prime}\big{[}\mathbb{P}\{A_{t}(\overline{a}_{t-1}^{\prime})=a_{t}\mid\overline{X}_{t}(\overline{a}_{t-1}^{\prime})\}\big{]}\mid X_{2},A_{1}=a_{1},X_{1}\right]w_{2}(\pi_{2})w_{2}^{\prime}(\pi_{2}^{\prime})\mid X_{1}\right)w_{1}(\pi_{1})w_{1}^{\prime}(\pi_{1}^{\prime})\right\}.$
@@ Appendix B Proofs for Section 5
$=\sum_{t=1}^{T}\int_{\overline{\mathcal{X}}_{t}}\widehat{m}_{t}\widehat{w}_{t}\mathbb{E}\left(\widehat{\phi}_{t}^{\prime}\mid\overline{a}_{t-1}^{\prime},\overline{x}_{t}\right)d\mathbb{P}(x_{t}\mid\overline{a}_{t-1},\overline{x}_{t-1})\prod_{s=1}^{t-1}w_{s}(\overline{x}_{s})w_{s}^{\prime}(\overline{x}_{s})d\mathbb{P}(x_{s}\mid\overline{a}_{s-1},\overline{x}_{s})$
@@
where the first line follows by repeating, from s=t1s=t-1 to $s=1$, iterated expectations on $\overline{X}_{s},\overline{A}_{s-1}=\overline{a}_{s-1}$ and then iterated expectations on $\overline{X}_{s-1},\overline{A}_{s-1}=\overline{a}_{s-1}^{\prime}$ and canceling terms, and the second line follows by repeating, from $s=1$ to t1t-1, iterated expectations on $\overline{X}_{s},\overline{A}_{s}=\overline{a}_{s-1}$.
@@
[…]
@@
Repeating this argument t2t-2 times, and noting that $\rho_{1}=1$, implies that $\mathbb{V}\{\varphi(Z)\}$ is bounded, which yields the result. ∎